LUCIFER Com 3, Night 10
UT 2009 January 19

Observers: N.Ageorges, W.Seifert, M.Juette, V.Knierim, J.Hill (@LBT)
Telescope Operator: D.Gonzalez-Huerta
SW Support: T. Edgin in Tucson
Telescope Support: T.Shih and J.Hill


The streak of spectacular good weather has ended. We started the night with thin cirrus, and the clouds got thicker as the night wore on. Closed for a couple hours starting at 02:25.

We sorted out the issues with global offsets on M1 and M2, so the collimation models now make a good starting point when the telescope is in temperature equilibrium.

Telescope was well-behaved until 6:30 AM when all kinds of strange TCS failures started happening. I can't tell if they are all related.


Butch worked alot on chiller CH0401 today, but he still hasn't solved the problem. His present suspicion is that the chiller isn't getting enough 45 degF water from from the utility building loop.

Stopped and restarted GCSL.

The Beginning of this Log has been preempted by the AFC Conference Championship Game

At the end of the third quarter we ran into the -Y limit on M1 again.

John confirmed that he was wrong about thinking he was wrong last night. M1 and M2 do want to move together with the same numerical sign of Xglobal and Yglobal. We reversed the sign of M2 in April 2008. John is still confused about why moving M1 and M2 in the same direction earlier in the week didn't help us get off the -Y limit, and we still don't really understand where the -5000 nm Z7 was coming (came) from. Thermal gradient in the telescope structure when the collimation model was last updated?

The new global offsets are:
M1: Xglobal=0 M2 Yglobal=-0.5 Zglobal=0.2 RX=+20 RY=0 (note that RX makes a pointing change also)
M2: Xglobal=-7.010 Yglobal=-0.200 Zglobal=0 RX=+114.28 RY=+181.80

IE=-31 CA=+10

Guide to coma seen on LUCIFER

Moving M1 positive gives coma like /\
Moving M1 negative gives coma like \/

Moving M2 positive gives coma like \/
Moving M2 negative gives coma like /\

So if they move the same direction the both coma and pointing are cancelled.

Moving RX positive gives coma like \/
Moving RX negative gives coma like /\

Sending positive Z7 gives coma like /\ (reversed compared to extrafocal LBC)
Sending negative Z7 gives coma like \/ (reversed compared to extrafocal LBC)

Moving RY positive gives coma like >
Moving RY negative gives coma like <

Ravens 14 -- Steelers 23 : making for an Arizona - Pittsburgh Super Bowl

LUCIFER is doing spectroscopy tests in very spectroscopic weather.

20:00 Sky is full of thin cirrus, with thicker stuff moving up from SW. T=+0.6 degC, D=-13 degC, wind (LBTO) 6 m/sec. Seeing varying from 1.2 to 2.0 arcsec.

20:37 put FCin=1 for Z11 and Z22.

20:43 Updated SXPMLUCIFERCollimation.dat and SXSMLUCIFERCollimation.dat to reflect the above values.

20:48 Updated PCSinstrument.conf and LUCIFERSX.20090118.ptmod so that the above global offsets are the defaults.

23:24 We still observing (testing acquisition software for spectroscopy), but the clouds are getting thicker (Polaris monitor can't see Polaris.) Seeing is about 1 arcsec.

02:25 Close for impending rain showers on the radar display.

Reopen (log is partially reconstructed since John was sleeping)

05:47 - Clouds have broken up, so we reopen.

06:29 preset ACTIVE - fails - Bad T0 polynomials

David restarted MCSPU and PCS.

~06:39 Active presets hanging. Got John out of bed. Killed all manner of susbsystems....PCS, GCSL, IIF. Finally found that it was MCS subsystem that was hung. Killing and restarting MCS freed up the preset.

07:00 preset ACTIVE - failed couldn't find star against background, and pointing was wrong.

Note the peculiar vignetting pattern on image acquire000118.

Reset IE=-33.4 CA=-7.95

07:02:03 preset TRACK - IIF thought this succeeded, BUT

7:02:38 Nancy got back a "no such handle error" from a StepFocus command that does NOT show on the IIFGUI. (neither the command nor the error showed up on the IIFGUI!)

I do find this in IIF syslog later: Jan 19 07:02:38 lbtmu102 LBT_IIF: Invalid handle to block: RPCERROR

07:03 preset TRACK - succeeds

07:07 Preset hangs again (looks like the same MCS symptom). Note that this command is still hung as of 07:28 AM (I'm leaving it in this stuck state.)

GCSL syslog shows: Jan 19 07:07:45 lbtmu107 LBT_left_GCS: Invalid handle in blockstub: �on�Xsn

07:10 Close dome for impending dawn.

07:24 At zenith in dark for LUCIFER calibrations

-- JohnHill - 17 Jan 2009
Topic revision: r5 - 19 Jan 2009, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback